Estlin Corporation sells mail-order computers. In December Year 2, it has generated $300,000 of sales revenue; the company expects a 20 percent increase in sales in January and 10 percent in February. All sales are on account. Estlin normally collects 80 percent of account receivable in the month of sale and 20 percent in the next month.

a. Prepare a sales budget for January and February Year 3

b. Determine the amount of sales revenue Estlin would report on the bimonthly pro forma income statement for January and February Year 3.

c. Prepare a cash receipts schedule for January and February Year 3.

d. Determine the amount of accounts receivable as of February 28, Year 3.
Answer #1

D C Sales Budget December January $ 300,000 $ 360,000 February 396,000 Sales $ m non Sales Revenue Janaury $ 360,000 Sales Fe

A D 1 C Sales Budget December January 300000 =B3+(B3* 20%) February =C3+(C3*10%) 3 Sales 5 February =D3 S ales Revenue Janaur

If you have any doubts ask in comments section

Do rate the answer